5t^2-10x=0
Can you show me exactly how 0 and 2 are possible answers to t.I know they are factors of 10 but I don't know how to set up the problem. Thanks

Answer by stanbon(75887)

5t^2-10t=0
Factor to get:
(5t)(t-2) = 0
Since the product is zero, one of the factors must be zero.
5t = 0 or t-2 = 0
Solve for "t" to get:
t = 0 or t = 2
====================
Cheers,
Stan H.